# GoalRank: Group-Relative Optimization for a Large Ranking Model

### v1

@@ Appendix C Reward Model Training
Appendix D Methods for Group Construction
@@
  • Autoregressive list generation (Jayaram & Thickstun, 2021): The conventional approach samples a single list by generating the entire trajectory in an autoregressive generator. While effective in capturing dependencies, this method is relatively slow and produces only one list per sampling trajectory.
  • Parallel tree-structured generation (Jayaram & Thickstun, 2021; Wang et al., 2025a): To improve efficiency and diversity, we allow the autoregressive generator to branch out in the first $K$ steps, forming a tree of partial sequences. The remaining LKL-K steps are then completed deterministically, enabling parallel exploration of multiple candidate lists.
  • Softmax-based stochastic sampling (Holtzman et al., 2019; Efraimidis & Spirakis, 2006): After the first step of scoring by generator, items are sampled probabilistically according to the softmax distribution of their scores, instead of deterministically selecting the top item, which encourages more diverse list generation.
@@
Appendix E Details of Offline Experiments Settings
E.1 Details of Dataset
@@ E.2 Details of Baselines
  • PRM (Pei et al., 2019a) addresses personalized re-ranking by integrating user-specific preferences into the re-ranking process, thus enhancing both personalization and relevance.

### v2

@@ Appendix B Reward Model Training
Appendix C Methods for Group Construction
@@
  • Autoregressive list generation (Jayaram and Thickstun, 2021): The conventional approach samples a single list by generating the entire trajectory in an autoregressive generator. While effective in capturing dependencies, this method is relatively slow and produces only one list per sampling trajectory.
  • Parallel tree-structured generation (Jayaram and Thickstun, 2021; Wang et al., 2025a): To improve efficiency and diversity, we allow the autoregressive generator to branch out in the first $K$ steps, forming a tree of partial sequences. The remaining LKL-K steps are then completed deterministically, enabling parallel exploration of multiple candidate lists.
  • Softmax-based stochastic sampling (Holtzman et al., 2019; Efraimidis and Spirakis, 2006): After the first step of scoring by generator, items are sampled probabilistically according to the softmax distribution of their scores, instead of deterministically selecting the top item, which encourages more diverse list generation.
@@
Appendix D Details of Offline Experiments Settings
D.1 Details of Dataset
@@ D.2 Details of Baselines
  • PRM (Pei et al., 2019b) addresses personalized re-ranking by integrating user-specific preferences into the re-ranking process, thus enhancing both personalization and relevance.
@@
To ensure a fair comparison, we adopt a relatively lightweight generator architecture for GoalRank. Specifically, the GoalRank Generator consists of two blocks: a lower block that performs feature crossing over all candidate items using several Transformer layers, and an upper block implemented as a Transformer decoder. During list generation, the decoder autoregressively predicts next-item scores over the full candidate set at each step, and GoalRank constructs the final list by sequentially selecting items (via Top-1 or sampling-based strategies).
D.3 Online Latency and MFU
To demonstrate the efficiency and resource-utilization advantages of GoalRank’s single-stage architecture, we report both the online latency and MFU of GoalRank compared with the existing Multi-Generator–Evaluator (MGE) pipeline. Notably, during training, many components of GoalRank can be executed in parallel. For example, the construction of auxiliary ranking-policy groups for reference-policy generation can be fully parallelized, resulting in negligible additional overhead. Under this setting:
  • Latency. GoalRank achieves an online latency of 18.611 ms, which is substantially faster than the multi-stage MGE pipeline (34.235 ms). This improvement stems from GoalRank’s ability to directly generate the final list in a single stage, eliminating evaluator scoring and multiple candidate-list constructions .
  • MFU. GoalRank attains an MFU of 12.65%, compared with 2.03% for the traditional two-stage MGE pipeline. The higher MFU reflects significantly better hardware utilization, ensuring that GoalRank does not increase overall training cost despite offering stronger performance.
These results validate that GoalRank is practical, efficient, and deployment-ready. Furthermore, GoalRank has been successfully deployed in our online environment to serve full user traffic.